A ball thrown vertically downward from the top of a building living the throwers hand with the velocity of 10 m per second. If the ball was through from a point 40 m above the ground how long does it take to strike the ground?

1
Expert's answer
2022-04-06T15:44:23-0400

"s=vt+\frac{gt^2}2,"

"t=\frac{\sqrt{v^2+2gs}-v}{g},"

"t=2~s."